# Analysis of reactor burnup simulation uncertainties for antineutrino spectrum prediction

### v2

@@ 3.1 Geometry and materials implementation
In order to simulate the fuel evolution in the SF95 and SF97 pins of the Takahama reactor, we implemented the geometry and the materials of the whole assembly in Serpent, taking as reference the first table in Takahama3 . Fig. 2 shows the assembly section in the xy plane of the simulated geometry. At the boundaries of the x and y directions, we apply reflective conditions to simulate the presence of the surrounding fuel assemblies, whereas in the z direction the geometry extends for 4.46 m, including the caps of the fuel cladding and a 20 cm thick water layer at top and bottom, to take into account the reflection of neutrons by water. To define the materials we used the cross sections of the ENDF/B-VII.1 ENDF_B_VII nuclear library.
3.2 Neutronic simulation of fresh fuel
Figure 3: Axial neutron flux of SF95 rod at the beginning of the depletion cycle.
Figure 4: Distribution of neutrons in the energy range in the SF95 rod location at the beginning of the depletion cycle.
The assembly was simulated at the beginning of the depletion cycle in the fresh fuel configuration, with the purpose of studying the spatial and energy distributions of neutron fluxes. In Fig. 3, we show the axial profile of the neutron flux intensity, obtained by setting an infinite energy domain in Eq. (5) and a spatial mesh with 100 equal subdivisions from $z=-182.4$ cm to $z=182.4$ cm, corresponding to the volume occupied by the fuel inside the pin. The bell shape of the neutron flux axial profile has the effect of burning the ${}^{235}$ atoms faster in the central part of the rod. As a consequence, as the fuel depletion proceeds, the axial profile becomes more flat, progressively equalizing the nuclide concentrations along the z axis. In Fig. 4, we show the neutron energy distribution, obtained by setting the spatial domain in Eq. (5) equal to the entire fuel volume and by dividing the energy range into 250 bins from $E=0.001$ eV to $E=20$ MeV (with the same width in logarithmic scale [^1] This subdivision of the energy bins is also called lethargic scale.).
@@ 3.4 Burnup simulation
We run Serpent simulations using an implicit scheme (Stochastic Implicit Euler, SIE) in order to avoid spatial instabilities that occur in 3D burnup calculations BurnupMethods . This method, in a similar way to all implicit schemes, solves the differential equation adopting the (unknown) end of step (EOS) flux value. At the beginning of each burnup step, a predictor passage of Forward Euler is applied. Then, several corrector iterations take place. Each iteration adopts the EOS flux from the previous corrector step. After repeating this procedure a fixed number of times, the final estimation of the nuclide density is obtained with the average of the corrector fluxes. More details about the SIE method can be found in SerpentSIE ; CastagnaImplicit ).
Figure 5: Geometries adopted for the description of fuel rods in the Takahama-3 reactor. (a) 30 burnup zones at the same temperature; (b) 30 axial burnup zones with 5 temperature zones for fuel and 3 temperature zones for water; (c) One burnup zone at uniform temperature. For consistency with respect to the experimental report the z-axis starts from the top of the fuel active region in the downward direction.
In order to map the evolution of nuclide concentrations along the axial direction and compare them with the experimental data, we divided the fuel pins into 30 burnup zones along the z axis. This means that in each volume the fuel burnup is computed starting from the neutron flux evaluated in it. Moreover, in order to estimate the effect of the temperature field characterizing the reactor core, we implemented three different simulations:
  1. (a)
    the same temperature $T_{\text{c,f}}=900$ K is applied to all 30 zones and water is set at $T_{\text{c,w}}=557$ K everywhere.
  2. (b)
    the 30 burnup zones are grouped into 5 macro-zones, applying respectively the temperatures of 750 K, 900 K, 1200 K, 900 K, 750 K. This profile is intended to reflect a realistic temperature distribution in the fuel, consistent with the shape of the neutron flux in Fig. 3, while preserving the average fuel temperature of 900 K. The water between the pins is divided in three zones with increasing temperatures from bottom to top (550 K, 557 K and 563 K).
  3. (c)
    a single burnup zone at uniform temperature is simulated for the whole assembly.
@@
The last case study (c) is not used for the purpose of the experimental benchmark due to the lack of spatial meshing, but is intended to investigate the role of spatial discretization in the computation of fission fractions. Fig. 5 shows a schematic of the three simulated models.
@@ 3.5 Results
Figure 6: Top panels: measured axial concentrations of the four fissile isotopes (black dots) in the SF95 pin, compared with simulation results obtained with uniform temperature field (blue) and different temperature zones (red). Bottom panels: C/E ratio of calculated (C) concentrations versus experimentally measured (E) ones for SF95 rod.
Figure 7: Top panels: measured axial concentrations of the four fissile isotopes (black dots) in the SF97 pin, compared with simulation results obtained with uniform temperature field (blue) and different temperature zones (red). Bottom panels: C/E ratio of calculated (C) concentrations versus experimentally measured (E) ones for SF97 rod.
The top panels of Figures 6 and 7 show the comparison between the measured axial concentration of the four fissile nuclides against the Serpent predictions. In these plots we also compare the simulation results obtained with the two models for the temperature field: (a) vs (b). Temperature acts directly on reaction rates and, thus, has an impact on fuel burnup. Particularly, at higher temperatures, the neutron flux Maxwellian distribution in the thermal range is peaked at higher energies where fission and capture cross sections are lower, but at the same time the Doppler broadening of cross section resonances has the effect of increasing the rates of reactions induced by neutrons in the intermediate range of the spectrum. Moreover, the temperature and, thus, the density of water has an impact on its moderation capability, which is more effective at lower temperatures. Depending on which of these effects is dominant, the reaction rates will be higher or lower as a function of the material temperature. For example, we observe that ${}^{239}$u accumulates more in the central part of the pin when this is simulated at a higher temperature. This is due to the fact that ${}^{239}$u is produced from neutron captures by ${}^{238}$, which mostly occur at the resonances.
In the bottom panels of Figures 6 and 7 we show the Calculated-to-Experimental (C/E) ratios, i.e. the ratios between the nuclide concentrations predicted by Serpent simulations and those measured experimentally. The closer the C/E ratios are to 1, the more accurate the prediction of nuclide concentration is. In principle, the uncertainty to be associated to C/E ratios is given by the propagation of both experimental and simulation uncertainties. The uncertainties reported in literature associated to the IDMS measurements are $<0.1\%$ for ${}^{235}$ and ${}^{238}$, and $<0.3\%$ for ${}^{239}$u and ${}^{241}$u. Since we run the Monte Carlo simulations with a relatively high number of neutrons (150 active cycles of 5×$\times 10^{10}$ neutrons/cycle), the statistical uncertainty of MC results is negligible with respect to that associated to the experimental data.
Looking at the C/E results shown in Figures 6 and 7, a very good agreement within ±$\pm 0.3\%$ is found, as expected, for ${}^{238}$, whose concentration remains very similar to the fresh fuel value. The other three fissile isotopes (${}^{235}$, ${}^{239}$u, and ${}^{241}$u) exhibit differences within the range $\pm$10in all sampling sites except for the uppermost ones (SF95-1, SF97-1, and SF97-2). By comparing the simulations performed with the two models of the temperature field (uniform vs axial profile), we observe local effects increasing or decreasing the nuclide concentrations up to 0.7%. Thus, the temperature field has a relatively small impact if compared with the systematic differences between experimental data and simulations.
@@
The data in the top locations show a worse agreement with the simulation results probably because these samplings were made close to the upper end of the fuel pin. In fact, the estimated uncertainty in the sampling position is $<2$ cm Suyama2008 . This is a negligible error in the central pin zone, where the flux profile becomes flat as the depletion goes on. On the contrary, at the pin edges where the neutron flux has a significant gradient, a 2 cm shift can introduce a non-negligible systematic error. Moreover, the neutronics at the axial edges depends on the back-scattering capability of the reflector. Unfortunately, the Takahama-3 reactor design data for reflector and cladding cap length and composition are not detailed in the technical reports. Their dimensions and compositions were hypothesised according to Suyama2008 , thus we simulated 2 cm of zircaloy-4 for cladding caps and a 20 cm layer of water.
For this reason, the following data analysis does not take into account the sampling performed in SF95-1, SF97-1 and SF97-2. Table 4 reports the average C/E for the considered samples (SF95 2-5 and SF97 3-6). In this table we also report, for comparison, the results published in literature obtained with other simulation tools (evaluating the average C/E avoiding the values at the rod extremes), namely: SWAT and ORIGEN 2 PIE_ORNL , SCALE 5.1 SCALE_PIE_ORNL and SCALE 4.4 and HELIOS Takahama3 . A similar analysis was previously conducted by Takahama_DRAGON_MURE , comparing a deterministic code (DRAGON) and a Monte Carlo (MURE).
The results obtained in this work with Serpent simulations are characterized by similar C/E values when compared with those obtained in other analyses. For what concerns the excluded data (i.e. SF95-1, and SF97 1-2), results from literature agree with our analysis, with differences with respect to unity up to 35%.
Rod Code ${}^{235}$ ${}^{238}$ ${}^{239}$u ${}^{241}$u
SF95 Serpent (a) 1.061 0.994 1.029 0.979
Serpent (b) 1.0422 1.0002 1.032 0.9577
@@
SF97 Serpent (a) 1.024 0.999 0.984 0.995
Serpent (b) 0.995 0.999 0.958 0.972

### v3

@@ 3.1 Geometry and materials implementation
In order to simulate the fuel evolution in the SF95 and SF97 pins of the Takahama reactor, we implemented the geometry and the materials of the whole assembly in Serpent, taking as reference the first table in Takahama3 . Fig. 2 shows the assembly section in the xy plane of the simulated geometry. At the boundaries of the x and y directions, we apply reflective conditions to simulate the presence of the surrounding fuel assemblies. In the z direction, the geometry extends for 4.46 m, including the caps of the fuel cladding and a 20 cm thick water layer at top and bottom, to take into account the reflection of neutrons by water, then, black boundary condition applies. Having the reflective boundary conditions in the x-y dimensions, this makes the assembly to burn as it was placed in the centre of the reactor core. During real cycle operations, assemblies are moved within the core structure when new fuel is added. Since no actual core position information were given, we acknowledge this extra uncertainty during the analysis. To define the materials in Serpent simulation, we used the cross sections of the ENDF/B-VII.1 ENDF_B_VII nuclear library.
@@ 3.3 Burnup simulation
The assemblies were simulated during the depletion cycle, starting from the fresh fuel configuration, with the purpose of studying the spatial and energy distributions of neutron fluxes. In Fig. 3, we show the axial profile of the neutron flux intensity at several depletion steps, obtained by integrating on the full energy domain in Eq. (5) and a spatial mesh with 100 equal subdivisions from $z=-182.4$ cm to $z=182.4$ cm, corresponding to the volume occupied by the fuel inside the pin. When the fuel is fresh, the neutron flux is spatially distributed with a bell shape, having the effect of burning the 235U atoms faster in the central part of the rod. As a consequence, as the fuel depletion proceeds, the axial profile becomes more flat, progressively equalizing the nuclide concentrations along the z axis.
Figure 3: Axial neutron flux of SF95 rod at several depletion steps, from the start of the cycle up to 800 days. Flux behaviour reflects the fuel evolution, progressively decreasing the number of neutrons in the central regions due to burnup. The zero of the axial position coincide with the top of the rod, accordingly to the Takahama report.
In Fig. 4, we show the neutron energy distribution at several depletion steps, obtained by setting the spatial domain in Eq. (5) equal to the entire fuel volume and by dividing the energy range into 500 bins from $E=0.001$ eV to $E=20$ MeV (with the same width in logarithmic scale [^2] This subdivision of the energy bins is also called lethargic scale.). In light water reactors, the energy region within $10^{-8}$ MeV and $10^{-6}$ MeV (also called, thermal energy region) plays a crucial role, since the fission interactions are more probable within this range. Depletion as well other design parameters, modify the evolution of the thermal region, having an impact on the reactor performances as well as on the fission fraction evaluation.
Figure 4: Energy distribution of neutrons within the SF95 fuel per unit lethargy. Fluence is represented for several depletion steps, from the start of the cycle up to 800 days. Flux behaviour within $10^{-8}$ MeV and $10^{-6}$ MeV is very important in light water reactors, being the region with the highest fission probability from fissile isotopes.
In order to map the evolution of nuclide concentrations along the axial direction and compare them with the experimental data, we divided the fuel pins into 30 burnup zones along the z axis. This means that in each volume the fuel burnup is computed starting from the neutron flux evaluated in it.
@@
We run Serpent simulations using an implicit scheme (Stochastic Implicit Euler, SIE) in order to avoid spatial instabilities that occur in 3D burnup calculations BurnupMethods . This method, in a similar way to all implicit schemes, solves the differential equation adopting the (unknown) end of step (EOS) flux value. At the beginning of each burnup step, a predictor passage of Forward Euler is applied. Then, several corrector iterations take place. Each iteration adopts the EOS flux from the previous corrector step. After repeating this procedure a fixed number of times, the final estimation of the nuclide density is obtained with the average of the corrector fluxes. More details about the SIE method can be found in SerpentSIE ; CastagnaImplicit ). The adopted statistic for each case is 160 active cycles + 100 inactive cycles in the predictor step, with 6×$\times 10^{5}$ neutrons/cycle and 16 substeps for the implicit scheme.
Figure 5: Description of the temperature and density profiles for the fuel and coolant along the z axis: solid horizontal lines are the values applied to the simulated Serpent regions, while the dashed lines are the continuous values based on design parameters. The zero of the axial position coincide with the top of the rod, accordingly to the Takahama report.
3.4 Temperature treatment
In order to estimate the effect of the temperature field treatment characterizing the fuel assembly, we implemented two different approaches:
  1. (i)
    the same temperature $T_{\text{c,f}}=900$ K is applied to all 30 zones [^3] The 30 zone divisions apply for all the fuel pins in the assembly. and water is set at $T_{\text{c,w}}=575$ K everywhere, with the correspondent density evaluated at p = 155.132 bar of $\rho_{\text{c,w}}=0.7227$ g/cm3. Average value data are applied coherently with the experiment report Takahama3 .
  2. (ii)
    the 30 burnup zones are grouped into 15 sub-zones, applying an axial dependent value for both the fuel temperature as well as the coolant temperature (and, consequently, density). Indeed, the coolant density value is evaluated from the combined knowledge of temperature/pressure with the pyXSteam pyXSteam library for Python. The coolant profile is chosen accordingly with the shape reported in the original report Takahama3 , constraining the inlet/outlet temperature to $T_{in}=557$ K and $T_{out}=594$ K. The fuel temperature profile is intended to reflect a realistic distribution, superimposing to the coolant temperature, a cosine-shaped curve, as reported in literature lamarsh2011introduction , constraining the average value to be equal to 900 K. Analytical expressions for coolant and temperature axial distributions are reported in Eq. (9)
    $T_{c}(z)$ $=T_{in}+c_{1}\left(1+\sin(c_{2}z+c_{3})\right)$ (9)
    $T_{f}(z)$ $=T_{c}(z)+d_{1}\cos(d_{2}z+d_{3})$
    The parameters {$c_{1},c_{2},c_{3},d_{1},d_{2},d_{3}$} are evaluated constraining inlet/outlet for coolant and average value for fuel. The analytical expressions along with the sub-zones averaged values adopted in the Serpent simulation are represented in Figure 5.
3.5 Boron treatment
[…]
  1. (i)
[…]
  2. (ii)
    Average boron treatment: the burnup calculation is carried out with a constant boron value, that we set equal to the mean value of the boron letdown curve discussed at the previous point.
Both approaches can be seen in Fig. 6, in which the amount of critical boron calculated by Serpent and its average (top panel, blue line and orange dashed line) to maintain the system’s reactivity equal to zero (bottom panel) is shown. Since we are simulating a single fuel assembly in an infinite lattice, starting from fresh fuel and without control rod insertion, we obtain a higher value of critical boron concentration with respect to the literature value (Takahama3 ). This analysis is able to check the difference among the results when using different boron treatments (constant boron vs variable boron). For sake of completeness, it is worth mentioning that an additional scenario is considered to complete the analysis, namely the burnup cycle with the imposition of the boron values reported in original report Takahama3 . Even though we acknowledge that this is theoretically the best approach for boron treatment, it is important to highlight that in current reactor operations, the boron value is not a priori known.
Figure 6: Evolution of the boron dissolved within the coolant according to the Serpent simulation: extra number of inactive cycles are added prior to the transport/depletion, evaluating the adjusted boron density (top panel) such that the system is critical (i.e., $\text{k}_{\text{eff}}$ is close to one). Bottom panel represents a measure of the system multiplication factor with the associated statistical uncertainty.
3.6 Simulated scenarios
Fig. 7 shows a schematic of the five simulated models, combining the temperature and boron treatments. The first three cases (A,B,C) are set in a constant boron environment. In particular,
@@ 3.6 Simulated scenarios
  • the first case study (A) is not used for the purpose of the experimental benchmark due to the lack of spatial meshing, but is intended to investigate the role of spatial discretization in the computation of fission fractions;
  • case (B) exploits a uniform temperature/density treatment, while case (C) imposes the spatially dependent properties;
  • cases (D) and (E) keep the same features of temperature treatment of (B) and (C) but with the automatic boron strategy.
The additional boron scenario mentioned in Section 3.5, only considered for validation purposes, adopts a uniform temperature profile and it will be labelled as (Ref). Temperature and density values applied to this schematic representation, are found in Fig. 5.
Figure 7: Geometries adopted for the description of fuel rods in the Takahama-3 reactor and for the simulation setups. (A) 1 burnup zone at constant temperature for fuel/coolant, in a constant boron environment; (B) 30 axial burnup zones at constant temperature for fuel/coolant in a constant boron environment; (C) 30 axial burnup zones with 15 temperatures for fuel/coolant in a constant boron environment; (D) 30 axial burnup zones at constant temperature for fuel/coolant in a variable boron environment; (E) 30 axial burnup zones with 15 temperatures for fuel/coolant in a variable boron environment; values for temperature in cases (C) and (E) is represented in Figure 5, and analytical expressions are found in Equation (9) . Constant boron values adopted for cases (A), (B) and (C) are evaluated as the integral mean from the boron letdown curve from cases (C) and (E). For consistency with respect to the experimental report the z-axis starts from the top of the fuel active region in the downward direction.
@@ 3.7 Results
This section shows the results of the validation between the Serpent model and the experimental data of the SF95 rod, for all the cases shown in Fig. 7. The purpose is to identify the modelling differences of a simulation by comparing it with the experimental data to quantify their importance in each case. The SF97 rod was compared with the simplest case (B), having the main purpose of testing the reliability of the model in the presence of an additional fuel cycle. Due to the minor importance for the purpose of this paper, the results of the SF97 rod are shown in Appendix A.
The top panels of Figures 8 show the comparison between the measured axial concentration of the four fissile nuclides against the Serpent predictions. In the bottom panels we show the Calculated-to-Experimental (C/E) ratios, i.e. the ratios between the nuclide concentrations predicted by Serpent simulations and those measured experimentally. The closer the C/E ratios are to 1, the more accurate the prediction of nuclide concentration is. In these plots we also compare the simulation results obtained with the four models for the temperature/boron field: (B), (C), (D), (E). As for the Uranium isotopes, both 235U and 238U appear to be almost insensitive to all setups: the former is reproduced with an average discrepancy from the experimental data of $\sim 12$%, while the latter remains below the $0.3\%$. The 239Pu turns out to be particularly sensitive to all the effects studied: in the spatial trend of this isotope, it is possible to identify and separate all the simulated case histories. In particular, the effect of constant boron vs. automatic boron turns out to be the predominant one: (D) and (E) results are lower (up to 10%) compared to their constant boron counterparts. The temperature approach, on the other hand, has a secondary effect of tilting the axial concentration, increasing it in the higher temperature zone. The effects are most visible on this isotope probably due to the fact that 239Pu is generated during burning, so its entire evolution is dependent on the boundary conditions used. The best case scenario is achieved with Simulation (E), which is able to catch the experimental data trend, with the lowest shift from them. The 241Pu shows similar trends to 239Pu, but less pronounced, and is better predicted.
In principle, the uncertainty to be associated to C/E ratios is given by the propagation of both experimental and simulation uncertainties. The uncertainties reported in literature associated to the IDMS measurements are $<0.1\%$ for 235U and 238U, and $<0.3\%$ for 239Pu and 241Pu. On the other hand, the uncertainty related to the simulation is not a trivial quantity to retrieve for what concern the results of the depletion scheme. The Serpent code is not able to directly estimate the error associated to the concentrations (C) coming from the solution of the Bateman equation, due to its deterministic formulation. For sake of simplicity, only average values of concentrations are considered, without a correlated uncertainty.
@@
It is worth mentioning that the data in the top locations (SF95-1, SF97-1 and SF97-2) show a worse agreement with the simulation results probably because these samplings were made close to the upper end of the fuel pin. In fact, the estimated uncertainty in the sampling position is declared $<2$ cm Suyama2008 . This is a negligible error in the central pin zone, where the flux profile becomes flat as the depletion goes on. On the contrary, at the pin edges where the neutron flux has a significant gradient, a 2 cm shift can introduce a non-negligible systematic error. Moreover, the neutronics at the axial edges depends on the back-scattering capability of the reflector. Unfortunately, the Takahama-3 reactor design data for reflector and cladding cap length and composition are not detailed in the technical reports. Their dimensions and compositions were hypothesised according to Suyama2008 , thus we simulated 2 cm of zircaloy-4 for cladding caps and a 20 cm layer of water.
For this reason, the following data analysis does not take into account the sampling performed in SF95-1, SF97-1 and SF97-2. Table 4 reports the average C/E for the considered samples (SF95 2-5). In Appendix A, the results for SF97 is reported for the samples SF97 3-6 in Table 6. A very good agreement is found for 238U whose concentration remains very similar to the fresh fuel value, with C/E falling within ±$\pm 0.5\%$. The other three fissile isotopes (235U, 239Pu, and 241Pu) exhibit differences depending on the sample site and the simulation case. However, an overall better agreement is found for the (E) case, exhibiting the best C/E for almost all sampling sites. By comparing the simulations performed with the two models of the temperature field (uniform vs axial profile), we observe local effects increasing or decreasing the nuclide concentrations up to 0.7%. Thus, the temperature field has a relatively small impact if compared with the systematic differences between experimental data and simulations. From the other side, the boron strategy is responsible for an higher difference, increasing the prediction performances up to 10% for the 239Pu. In these tables we also report, for comparison, the results published in literature obtained with other simulation tools (evaluating the average C/E avoiding the values at the rod extremes), namely: SWAT and ORIGEN 2 PIE_ORNL , SCALE 5.1 SCALE_PIE_ORNL and SCALE 4.4 and HELIOS Takahama3 . A similar analysis was previously conducted by Takahama_DRAGON_MURE , comparing a deterministic code (DRAGON) and a Monte Carlo (MURE). The results obtained in this work with Serpent simulations are characterized by similar C/E values when compared with those obtained in other analyses. For what concerns the excluded data (i.e. SF95-1, and SF97 1-2), results from literature agree with our analysis, with differences with respect to unity up to 35%.
Figure 8: Top panels: measured axial concentrations of the four fissile isotopes (black dots) in the SF95 pin, compared with simulation results obtained with the design setups B, C, D and E, previously defined. Bottom panels: C/E ratio of calculated (C) concentrations versus experimentally measured (E) ones for SF95 rod.
Rod Code 235U 238U 239Pu 241Pu
SF95 Serpent (B) 1.124 1.002 1.164 1.058
Serpent (C) 1.120 1.001 1.175 1.07
Serpent (D) 1.110 1.003 1.081 0.991
Serpent (E) 1.111 1.002 1.087 0.997
Serpent (Ref) 1.083 1.003 1.042 0.955
@@
Table 4: Average Serpent C/E for chosen samples (SF95 2-5) in the selected cases B, C, D, E and Ref, compared with the literature results obtained with different simulation codes.
4 Fission fractions and $\bar{\nu}_{e}$ spectra results
4.1 Fission Fractions
Among the possible combinations of cases simulated according to the scheme in Fig. 7, we want to give importance to the principal effect, which was also shown in the validation reported in the previous section, namely, the choice of the boron strategy. Figure 9 shows the trend of fission fractions in the standard fuel in the case with uniform temperature and constant boron (B) versus uniform temperature and automatic boron (D), with their relative difference. With fresh fuel, the contribution to the fission events comes from 235U ($\sim$ 94%) and 238U ($\sim$ 6%). As the fuel depletion proceeds, two major effects can be appreciated: 235U is consumed and Pu nuclides are produced from 238U neutron captures. Therefore, with the increase in the burnup level, the relative importance to the total fission events of 235U decreases, while the 239Pu and 241Pu contribution increases. The different background colors represent the two simulated cycles, each characterized by a different nominal power. For the presented case, the (B)-(D) comparison shows a clear dependence on the burnup level: the relative difference between the fraction increases as a function of the burnup level. In particular, the fraction of the 238U is the one most sensible to the boron treatment. To explain this phenomenon, we should look at what happens in the thermal region of the neutron spectrum in the different cases: as depicted in Fig. 10, the change in the boron level consist in a very large modification of the Maxwellian peak: when the boron concentrarion is lower, the peak is higher and viceversa. This fact influences how many fissions occur in fissile isotopes (235U, 239Pu or 241Pu) and how many in the fissionable (238U).
Figure 9: Comparison between fission fraction in simulation B (green line) and simulation D (orange line) and their relative difference (red dashed line). The effect of the boron treatment (automatic against average) on the fission fraction calculation shows a trend for every fission fraction. Fission events in 238U are the one that suffer most from the boron level in the system, completely changing the slope of the time evolution, with a maximum difference of -10 %.
Figure 10: Neutron fluence in the thermal energy range for the various simulation cases (B to E) during the burnup evolution (from 0.0 MWd/kg up to 30.74 MWd/kg). Simulation setups where automatic boron is activated (i.e., D and E) shows an higher change in the neutron energy distribution in the range where fission is more probable (for 235U, 239Pu and 241Pu).
Table 5 report the cycle-averaged fission fractions for all the presented scenarios, for both cycle 5 and cycle 6. Again, the highest differences can be appreciated comparing automatic boron cases against constant boron cases. Here, is reported the relative difference in the (B)-(D) comparison for both cycles:
  • (235U) $-0.5\%$ for cycle 5 and $+2.0\%$ for cycle 6;
  • (238U) $+3.5\%$ for cycle 5 and $-5.4\%$ for cycle 6;
  • (239Pu) $+0.5\%$ for cycle 5 and $-1.5\%$ for cycle 6;
  • (241Pu) $+0.2\%$ for cycle 5 and $-1.1\%$ for cycle 6.
Another interesting difference can be appreciated comparing the cycle averaged fractions of case (A) against the others: the effect of the number of burnup zones is much more pronounced in the first cycle with respect to the second. The precise comparison of the other cases (temperature effect, mesh effect) will be faced in Appendix B. For what concern the boron simulation strategy, we can state that its modelling in MC simulations has a non-negligible effect in the estimation of fission fractions, especially the 238U one, where a non correct approach may lead to a fission fraction overestimation/underestimation, depending on the cycle.
Case 235U 238U 239Pu 241Pu
Cycle 5 (A) 0.734 0.068 0.185 0.012
(B) 0.724 0.069 0.193 0.014
(C) 0.725 0.069 0.192 0.014
(D) 0.721 0.071 0.194 0.014
(E) 0.720 0.072 0.194 0.014
Cycle 6 (A) 0.502 0.078 0.359 0.061
(B) 0.502 0.078 0.358 0.062
(C) 0.500 0.079 0.359 0.062
(D) 0.512 0.074 0.352 0.061
(E) 0.510 0.075 0.354 0.061
Table 5: Cycle-averaged fission fraction for the simulated scenarios.
4.2 Non oscillated $\bar{\nu}_{e}$ spectra results
Figure 11: Top: $\bar{\nu}_{e}$ spectra multiplied by the IBD cross section computed with the fission fractions corresponding to the cycle averages value in the simulation (B). Bottom: Relative differences of the two spectra.
Figure 12: Top: $\bar{\nu}_{e}$ spectra multiplied by the IBD cross section computed with the last averaged cycle fission fractions for the cases considered. Bottom: Relative differences of the spectra with respect to the (A) case.
We finally present the equilibrium $\bar{\nu}_{e}$ spectra obtained through the summation method by combining the cumulative fission yield (CFY) data published in the Live Chart of Nuclides by IAEA IAEA_livechart and the $\bar{\nu}_{e}$ spectra of 468 $\beta^{-}$ decays obtained through BetaShape by setting an energy binning of 1 keV. We use thermal CFY for 235U, 239Pu and 241Pu, which are mostly fissioned by thermal neutrons, and fast CFY for 238U which can be fissioned with not negligible probability by fast neutrons only. Then, the total reactor $\bar{\nu}_{e}$ spectrum (Equation (2)) is evaluated with the fission fractions estimated with Serpent simulations. Since the reactor $\bar{\nu}_{e}$ will be detected through the IBD reaction, we multiply the $\bar{\nu}_{e}$ spectra by the IBD cross section STRUMIA200342 so as to obtain the shape of the detectable non oscillated spectrum. We want to highlight that this analysis is meant to compare the propagation of the reactor model differences on the reactor antineutrino spectrum. Other reactor antineutrino calculation methods as well as comparison with experimental data is out of the scope of the analysis and will be presented in future works.
In Fig. 11, we show the $\bar{\nu}_{e}$ spectra obtained by using the cycle-averaged fission fractions estimated by Serpent for the Takahama-3 assembly, highlighting the differences between the emission spectrum during the first cycle and the second one. As shown in the top pad, the $\bar{\nu}_{e}$ flux intensity decreases with time. This is because the average number of $\bar{\nu}_{e}$ emitted after 239Pu fission is lower with respect to that of 235U. The depicted error band for each spectrum propagates the uncertainties of the following variables, namely Q-values, branching ratios and fission yields. For sake of simplicity, the uncertainties are treated as independent. The bottom pad shows the relative difference calculated with respect to the fresh fuel $\bar{\nu}_{e}$ flux is of the order of -5% around the maximum of the spectrum at 4 MeV, and is up to $\sim-10\%$ in the high energy range. This result reflects the extreme case where the reactor starts from the fresh condition, without any refuelling strategy: in a more realistic scenario, this effects is averaged out considering fuel at different depletion stages. Nevertheless, in the framework of the JUNO experiment, it will be very important to correctly sum up the non-oscillated spectra taking into account the fission fraction evolution of each of the eight reactors during the data taking.
Finally, in Fig. 12 we show the impact of the different simulated scenarios discussed in Sect. 4.1 on the non oscillated $\bar{\nu}_{e}$ spectrum. In particular, the spectra shown in the top pad were computed by using the cycle-averaged fission fractions for the second cycle, obtained from the simulations (A), (B), (C), (D) and (E). In this case, the differences are much less pronounced, remaining below 0.2% up to 7 MeV and in any case below 2% at higher energies. This can be addressed to the fact that the major differences are related to the 238U fraction, which has a minor impact on the overall system.
5 Conclusions
In this paper we presented a methodology for the prediction of the non-oscillated reactor $\bar{\nu}_{e}$ spectrum. This is intended to serve the scientific community in future reactor $\bar{\nu}_{e}$ experimental campaigns (e.g., JUNO and TAO) that will analyze the mass ordering problem with an unprecedented level of precision. We coupled the Monte Carlo code Serpent along with the $\bar{\nu}_{e}$ spectrum calculation software BetaShape. The Serpent simulations allow the calculation of fission fraction evolution over time, leveraging data encompassing reactor geometry, power history, and fresh fuel inventory. Subsequently, by integrating BetaShape alongside nuclear databases, we construct the antineutrino spectra emanating from the four primary fissile isotopes with the summation method – 235U, 238U, 239Pu, 241Pu – thereby enabling predictions of the total reactor antineutrino spectra across various burnup levels.
In this paper we focused our analysis on the fuel burnup, aiming at exploring the accuracy in the evaluation of fission fractions through MC simulations. For this purpose, we used the spent fuel concentrations of the Takahama-3 PWR as experimental benchmark for the Serpent simulations. We found that Serpent simulation run with either uniform temperatures in the fuel/moderator or imposing different temperature zones, are both in a reasonably good agreement with the experimental data. Importantly, our simulations highlight that the temperature field introduces minor discrepancies both in the fuel depletion calculations as well as when considering the fission fractions. We also focused on the boron strategy, comparing the usage of an automatic critical boron calculator and the approach with an constant boron density. When the automatic critical boron is present, the thermal neutron flux is highly influenced by this field, leading to a not-negligible difference in the concentration estimation (C/E improved by 10%) as well as in the fission fraction evaluation (relative difference up to 5% ).
Looking ahead, our future plans encompass harnessing this analytical tool to further investigate additional sources of systematic uncertainty (such as cross-section or fission yield data uncertainties) and their subsequent impact on the antineutrino spectrum. Future developments will be devoted to relax the equilibrium hypothesis on the fission products, highlighting the importance that this amount of nuclide have on the spectrum. Also, a more detailed analysis of the reactor configuration of the Taishan and Yangjiang plant will be addressed, along with the comparison with other available experimental spectra from near-detector experiments.
6 Data availability
The authors confirm that the data supporting the findings of this study are available within the article.
Appendix A
Appendix A SF97 validation
This appendix shows results for the SF97 rod validation, being analyzed in the simplest case, considering a uniform temperature for fuel and coolant, and a constant boron approach (i.e., simulation (B)).
Figure 13: Top panels: measured axial concentrations of the four fissile isotopes (black dots) in the SF97 pin, compared with simulation results obtained with uniform temperature field and average boron approach. Bottom panels: C/E ratio of calculated (C) concentrations versus experimentally measured (E) ones for SF97 rod.
The top panels of Figures 13 show the comparison between the measured axial concentration of the four fissile nuclides against the Serpent predictions, in the scenario (B). Bottom panels show the C/E ratio, giving a quantitavive measure of how the simulation is close to the measurement. Overall, the comparison shows a good agreement, despite the first sample SF97-1, which is overestimated/underestimated up to a factor 2 for 241Pu. Problems about comparison with the first two samples are reported in Sect. 3.7. Finally, Table 6 shows the average C/E analysis compared with other literature references.
Rod Code 235U 238U 239Pu 241Pu
SF97 Serpent (B) 1.040 1.002 1.078 1.038